OpenStudy (anonymous):

How many units of MgSO4 are present in 18.0g of this compound? units as in mass? molecules? or atoms?

OpenStudy (abb0t):

Use dimensional analysis to convert to moles. This can be see by the relationship on the periodic table - molar mass. Which means their units are: \(\frac{ g }{ mol }\). Then, similar to how you did in algebra, cancel units. Then use avogadro's number to convert to molecules/ atoms: \(6.022 \times 10^{23} \frac{ molecules/atoms }{ mol }\)

OpenStudy (anonymous):

so by the word "units" it's asking for grams?

OpenStudy (abb0t):

Yes. Units is some type of measurement.

OpenStudy (anonymous):

9.00 x 10^22?